OpenStudy (anonymous):

9x^4-21x^2+10=0 solve!

OpenStudy (jlastino):

(3x^2-2)(3x^2-5) x= sqrt (2/3) x=sqrt (5/3)

OpenStudy (saifoo.khan):

wrong!!

OpenStudy (saifoo.khan):

factors are needed.

OpenStudy (anonymous):

so how would i write it?

hero (hero):

9x^4-15x^2-6x^2+10 = 0 3x^2(3x^2-5)-2(3x^2-5) = 0 (3x^2-5)(3x^2-2) = 0 x = +/- sqrt{5/3} x = +/-sqrt{2/3}

OpenStudy (saifoo.khan):

Still NO.

hero (hero):

Well Legend....Since you seem to know the solution, why don't you reveal it oh great one...

OpenStudy (saifoo.khan):

Well Zero, factors are needed. I know

hero (hero):

I guess wolfram alpha is wrong too

hero (hero):

Now I get to call you Absolute Zero

OpenStudy (saifoo.khan):

lol, Molly wrote the question incorrect.

hero (hero):

No, no....she didn't. The only angle of incorrectness came spewing from your big mouth which never ceases to shut itself.

OpenStudy (saifoo.khan):

whtever you say, but i know that the question she wrote incorrect. solutions aren't needed in that question but facrots are. :p

OpenStudy (saifoo.khan):

factors*

hero (hero):

Whatever. I guess you have a copy of her homework notes her teacher gave her too. I guess you can see through computer screens, read minds, and walk through walls. Oh great one, why did I ever doubt you?

OpenStudy (saifoo.khan):

haha, infact i do have a copy of those questions! :P

OpenStudy (anonymous):

i dont think im going to finish the exam in time !!!! :(

OpenStudy (anonymous):

ahhhh

hero (hero):

Okay. I guess you must be taking the same online test she is. Either way, you can't say it's wrong if we correctly posted the solution for the question she did write. Go away Absolute Zero.

OpenStudy (saifoo.khan):

I never said that solutions are wrong. Solutions are right, but the answers in which form they are needed, they are incorrect.

hero (hero):

Let me guess....another U of P online tester....

hero (hero):

U of P obviously got their math program from the Illuminati wing of the Mafia.

OpenStudy (saifoo.khan):

Nah, not only UoP.. Others use it as well.

hero (hero):

Yep, and they all signed contracts with the Illuminati and Mafia

OpenStudy (saifoo.khan):

Agree. ^_^